We find max and min value of


asin@+bcos@ by +-( underroot ( a power2 + b power2))


Can we find the max and min value of asin(2@)+bcos(2@) using the same formula as +-( underroot ( a power2 + b power2)).


Please explain in detail. Urgently..

Grade:12

1 Answers

Jaya IITK
12 years ago

Get into the depth why u use that formula.

Its very important for iit point of view.

c=a sin@ + b Cos@

c/under root(a^2+b^2)= (a /root under(a^2+b^2)) sin@+ (b/root under(a^2+b^2)                                                                                                             cos@

 

If u take a right angled triangle witrh sides a, b then its hypotenues will be under root(a^2+b^2)

 

 

therefore its in the form of:

 

c/root(a^2+b^2) = cos $ sin @ + sin$ cos @

 

c= root(a^2+ b^2) sin($+@)

 

-1<=sin(@+$)<= 1

 

therefore c lies between - root( a^2+b^2) and root(a^2+ b^2)

 

thus u can see even if u replace @ by 2@ u get the same result..

 

Plzzzzzzzzzzz approve my answer!!!!!!!!!